OpenStudy (ajprincess):

Please help:) Show the state of the stack and the value of each variable after execution of each of the following statements: p=2.5 q=3.30 r=7.003 d=0.00026 create stack push q onto stack push d onto stack pop item from stack push r onto stack push p onto stack p=q*r push p+r+d onto stack pop item from stack and store in p pop item from stack and store in q pop item from stack and store in r

OpenStudy (tyteen4a03):

Create Stack: *bzzt* Push q onto stack, stack becomes [q] Push d onto stack, stack becomes [q, d] pop item from stack, stack becomes [d] push r onto stack, stack becomes [d, r] push p onto stack, stack becomes [d. r. p] p=q*r, stack becomes [d, r, q*r] push p+r+d onto stack, stack becomes [d, r, q*r, p+r+d] pop item from stack and store in p, p becomes d pop item from stack and store in q. q becomes r pop item from stack and store in r, r becomes q*r.

OpenStudy (konradzuse):

^That is wrong.

OpenStudy (konradzuse):

create stack push q onto stack[q] push d onto stack[q,d] pop item from stack[q] push r onto stack[q,r] push p onto stack[q,r,p] p=q*r // do nothing push p+r+d onto stack[q, r, p, (p[q * r]+r+d)] pop item from stack and store in p [q,r,p] // (p[q * r]+r+d)] = p pop item from stack and store in q [q,r] // p = q pop item from stack and store in r [q] // r = r

OpenStudy (ajprincess):

@KonradZuse Can u please explain me the line push p+r+d onto stack[q, r, p, (p[q * r]+r+d)] Why it is p[q*r]? I dnt get that part.

OpenStudy (konradzuse):

I meant p = q * r, syntax fail on my part.

OpenStudy (konradzuse):

p + r + d or (r*d) + d + r

OpenStudy (ajprincess):

u mean p+r+d=(q*r)+d+r?

OpenStudy (ajprincess):

@KonradZuse so it will be create stack push q onto stack[q] push d onto stack[q,d] pop item from stack[q] push r onto stack[q,r] push p onto stack[q,r,p] p=q*r // do nothing push p+r+d onto stack[q, r, p, ([q * r]+r+d)] pop item from stack and store in p [q,r,p] // ([q * r]+r+d)] = p pop item from stack and store in q [q,r] // p = q pop item from stack and store in r [q] // r = r. Isn't t?

OpenStudy (tyteen4a03):

@KonradZuse D'oh, messed up stacks with queues.

OpenStudy (konradzuse):

yes remember queues are first in first out(think of a line of school children, or lunch line) then stacks are first in last out, so think of a stack of papers. If you put a papper at the bottom, you need to do all the papers on top of it first to get to that one paper.

OpenStudy (konradzuse):

not too sure why r = r, but that's how it's supposed to be done.....

OpenStudy (ajprincess):

oh k. thank u soooo much @KonradZuse

OpenStudy (konradzuse):

np, goodluck.